OpenStudy (liizzyliizz):

given the curve 4x^2+2xy-xy^3=3 find the values of dy/dx at the point (1,-1)

OpenStudy (bahrom7893):

implicit differentiation

OpenStudy (anonymous):

\[4x^2+2xy-xy^3=3\] \[8x+2xy'+2y-y^3-3xy^2y'=0\] replace x and y by the numbers you know and solve for \[y'\]

OpenStudy (bahrom7893):

8x + 2 * (x * dy/dx + y) - (x * 3y^2 * dy/dx + y^3) = 0

OpenStudy (bahrom7893):

Ughh I keep crashing for some reason... now just do whatever satellite said.

OpenStudy (liizzyliizz):

So would it be 7? I am not sure I probably messed something up. I will double check.

OpenStudy (anonymous):

now it is arithmetic i can try it, maybe

OpenStudy (anonymous):

\[8x+2xy'+2y-y^3-3xy^2y'=0\] \[8+2y'-2-1-3y'=0\]

OpenStudy (anonymous):

i replaced each x by 1 and each y by -1, but i am not guarenteeing my answer

OpenStudy (anonymous):

this gives \[5-y'=0\] or \[y'=5\] again, i would not bet my lunch money on it, but it looks right

OpenStudy (liizzyliizz):

ohh i added the one in stead of subtracting :l ugh the dangers of negative signs lol